OpenStudy (anonymous):

Indicate the x- and y-intercepts. ( Equation: f(x) = 4x^3 + 6x^2 − 24x + 1 So I set it=0 but I don't know how to solve this :/ please help! 4x^3 + 6x^2 − 24x + 1=0

OpenStudy (anonymous):

Are you in calculus?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. Getting the first derivative, y' = 12x^2 + 12x - 24 = 12(x + 2)(x - 1), we have extrema at x= 1 or -2. With me so far?

OpenStudy (anonymous):

yes, but I thought all you had to do for finding the x-intercept was to set the equation=0?

OpenStudy (anonymous):

Yes, that's correct, and then you use synthetic division to look for rational zeros and also use Descartes' Rule of signs. But there are no rational zeros, just irrational but real zeros. I was going into the derivative so that you could roughly graph it so that you see the curve, but we don't have to.

OpenStudy (anonymous):

oh I see, extrema would the y value just be 0? so -2,0? 1,0?

OpenStudy (anonymous):

 Well, the extrema are (-2, 9) and (1, -13). You have a y-intercept at (0, 1). You have an irrational zero at slightly less than x = 1/4.